# Black-box Adversaries from Latent Space: Unnoticeable Attacks on Human Pose and Shape Estimation

### v1

@@ 4.2 The proposed Method
where $t$ denotes the iteration steps, $\hat{z}^{(t)}$ represents the latent perturbation at step $t$, and $\xi_{z}$ is the learning rate for the gradient update in latent space. The term $\frac{\partial\mathcal{G}_{\theta}}{\partial\hat{z}^{(t)}}$ denotes the Jacobian matrix of the Decoder, whose computation is prohibitively expensive due to its high dimensionality ($n\times d$, often very large). Furthermore, each update necessitates a complete forward pass through the decoder network, resulting in substantial memory and computational consumption. Consequently, performing iterative updates directly in the latent space is highly inefficient (see Method C in Table 3).
$\Delta x^{(0)}=x^{(0)}-x,\quad\Delta x^{(0)}=\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2}).$ (16)
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{x}\cdot\nabla_{\Delta x^{(t)}}\mathcal{L}(\mathcal{P}(x+\Delta x^{(t)}),(\alpha,\beta,\gamma)),$ (17)
$\nabla_{\Delta x^{(t)}}\mathcal{L}(\mathcal{P}(x+\Delta x^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial(x+\Delta x^{(t)})},$ (18)
where $\xi_{x}$ denotes the learning rate for gradient updates in pixel space. Importantly, computing gradients in pixel space does not require backpropagation through the decoder network. Instead, perturbations are applied directly to the original image $x$, thereby eliminating the need to compute the high-dimensional Jacobian transformation from latent to pixel space. This approach substantially reduces both the computational complexity and memory usage of each gradient update step.
$\mathcal{L}=-\mathbb{E}_{(\hat{\alpha},\alpha,\hat{\beta},\beta,\hat{\gamma},\gamma)}\left[\|\hat{\alpha}-\alpha\|^{2}+\|\hat{\beta}-\beta\|^{2}+\|\hat{\gamma}-\gamma\|^{2}\right]+\frac{1}{w\times h}\sum_{i=0}^{w-1}\sum_{j=0}^{h-1}\left\|\hat{x}_{i,j}-x_{i,j}\right\|^{2},$ (19)
where $w,h$ denotes the width and height of $x$. The first term enforces the perturbed pose parameters to deviate from the ground-truth parameters (enhancing the attack effectiveness), while the second term strictly constrains pixel-level variations to maintain the imperceptibility of the perturbation. According the Eqs 17, 18 and 19, the update rule for the pixel-space perturbation is formulated as:
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{x}\cdot(\frac{\partial\mathcal{L}}{\partial\mathcal{P}}\frac{\partial\mathcal{P}}{\partial(x+\Delta x^{(t)})}-\lambda\cdot(x^{(t)}-x)),$ (20)
where $\lambda$ is a hyperparameter that balances attack strength and imperceptibility. It is evident that the computational complexity of updates in latent space is $\mathcal{O}(n\times d)$, whereas the complexity of pixel-space updates is only $\mathcal{O}(n)$. Therefore, the proposed noise enhancement strategy in pixel space incurs significantly lower computational overhead (see the Method C and UBA (Ours) in Table 3).
Notably, we interact with the model solely via its application interface (API), without access to internal parameters, and impose a strict limit on the number of queries $t$. This design reduces the risk of detection due to excessive querying and alleviates the burden on the attacker.
5 Experiments
5.1 Experimental Settings
Datasets and EHPS Models. We evaluate our method on two benchmarks: 3DPW [38], with in-the-wild RGB videos and 3D SMPL pose annotations, and UBody [24], a large-scale upper-body mesh dataset with 1.05 million annotated images across 15 scenarios. For black-box evaluation, we test UBA on state-of-the-art EHPS models, including Hand4Whole [28], OSX [24], and SMPLer-X [2], which provides four ViT-based pre-trained variants, denoted as SMPLer-X-M, where M denotes the size of the Vision Transformer (ViT) backbone [1] (M $\in$ S, B, L, H).
Baseline and Evaluation Metrics. Currently, only one adversarial attack targets EHPS. To evaluate UBA, we compare it with state-of-the-art methods in computer vision, including classical attacks (FGSM [8], PGD [27]), diffusion-based approaches (ACA [5], DiffAttack [4]), and the EHPS-specific TBA [23]. We report MPJPE and MPVPE for 3D joint and mesh accuracy, and PA-MPJPE/PA-MPVPE for alignment-aware evaluation, following prior EHPS benchmarks [28, 24, 2]. All metrics are in millimeters (mm).
Implementation Details. The experiments are conducted on an Ubuntu 20.04.6 LTS system using the PyTorch framework, a single NVIDIA L20 GPU, and 45 GB of RAM. The VAE model employed is stabilityai/sd-vae-ft-mse from Hugging Face  [^1] An open source platform. https://huggingface.co/stabilityai/sd-vae-ft-mse. We set the number of queries to $t=3$, the perturbation magnitude to $\eta=0.05$, and the balancing hyperparameter to $\lambda=0.5$.
5.2 Attack Performance on EHPS Models
@@
Table 1: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm)
@@
UBA (Ours) 56.59 (37.89%) 9.74 (24.08%) 3.52 (18.92%) 128.57 (37.05%) 54.02 (51.74%) 41.83 (38.05%) 71.67 (43.92%) 9.93 (24.59%)
Figure 2: Visualizing various adversarial samples for digital human generation.
We first evaluate the effectiveness of UBA on the UBody dataset, which includes full-body, hand, and facial regions of digital humans. As shown in Table 1, across all tasks, the adversarial examples generated by UBA cause the most significant degradation in the accuracy of all EHPS models, compared to other attack methods. Specifically, the average increase in estimation error ranged from 25.13% to 58.21%. Notably, for the SMPLer-X-H model, the MPVPE (All) increased from 41.22 mm to 88.60 mm, corresponding to a 114.94% error growth rate.
@@
Table 2: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the 3DPW dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack MPJPE (Body) $\downarrow$ (mm) PA MPJPE (Body) $\downarrow$ (mm)
@@
UBA (Ours) 128.72 (27.88%) 80.10 (17.92%)
(a) UBody
(b) PW3D
Figure 3: Attack performance of different perturbation magnitude $\eta$.
(a) UBody
(b) PW3D
Figure 4: Attack performance of different query times $t$.
provided in the Appendix B). As shown in Fig. 4, the number of queries $t$ is positively correlated with the mean error growth rate, suggesting that increasing the number of queries enhances attack effectiveness. Notably, when $t=9$, the mean error growth rates for SMPLer-X-B, SMPLer-X-L, and SMPLer-X-H all exceed 100%. However, an excessive number of queries raises the attacker’s
We further evaluate UBA on the 3DPW dataset, focusing on how adversarial examples impact the estimation of body parts across different EHPS models. As presented in Table 2, UBA consistently causes the greatest reduction in model accuracy. In particular, UBA yielded a minimum error growth rate of 17.92% (increasing PA MPJPE from 67.93 mm to 80.10 mm on the Hand4Whole model) and a maximum of 48.46% (raising MPJPE from 79.46 mm to 118.00 mm on the SMPLer-X-B model), with average increases ranging from 17.27% to 44.49%.
These results demonstrate that UBA is highly effective and destructive. Remarkably, it functions as a black-box attack, requiring minimal prior knowledge—achieving success through API queries alone. Interestingly, we observe negative error growth rates in a few instances, such as PA MPVPE (Face) under the OSX model on UBody, and several other metrics under specific settings, suggesting that adversarial attacks may not only impair EHPS performance but could, in rare cases, enhance estimation accuracy.
To further illustrate the effectiveness of UBA, we conduct qualitative evaluations by visualizing the digital humans generated from both clean and adversarial samples produced by various attack methods. As shown in Fig. 2, UBA leads to noticeable degradation in pose and shape estimation accuracy, resulting in significant deviations in the synthesized digital humans. These visual results clearly highlight the adverse impact of UBA on the EHPS models’ ability to accurately reconstruct human poses (Please refer to Appendix B for more results).
Table 3: Ablation study of different adversarial attack frameworks on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Method PA MPVPE $\downarrow$ (mm) MPVPE $\downarrow$ (mm) PA MPJPE $\downarrow$ (mm) Memory
@@ C Out of Memory $\geq$ 45 GB
UBA (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%) 7.47 GB
5.3 Ablation Study
@@
Effects of Different Noise Injection Strategies. To further explore the advantages of UBA, we revisit three alternative and more direct noise injection strategies within the UBA framework: A) adding random noise in the pixel space and iteratively updating it by querying the EHPS model; B) injecting random noise into the latent space after VAE encoding, followed by resampling and decoding to generate adversarial examples; and C) injecting random noise into the VAE latent space, then resampling and decoding, with iterative updating it by querying the EHPS model. We evaluate these strategies on the SMPLer-X model using the UBody dataset and report the GPU memory required for generating adversarial examples. As presented in Table 3, our method achieves strong attack performance while requiring only 7.47GB–12.83GB of GPU memory, indicating high practicality. These results demonstrate the effectiveness and efficiency of our approach in attacking EHPS models without the need for extensive computational resources.
@@
Effects of Different Perturbation Magnitude $\eta$. To gain a clearer understanding of the impact of varying perturbation magnitudes $\eta$ on the EHPS model, we conduct experiments using the SMPLer-X model and report the mean error growth rates across all tasks on the UBody and PW3D datasets, respectively (task-specific results are provided in the Appendix B). As illustrated in Fig. 3, different perturbation magnitudes $\eta$ result in comparable attack effectiveness. This phenomenon may be attributed to the model’s sensitivity to latent-space perturbations being concentrated along specific directions rather than being strictly dependent on the perturbation magnitude. As a result, increasing the magnitude does not substantially improve attack performance.
Effects of Different Query Times $t$. Similarly, we conduct an ablation study on the number of queries to the EHPS model API, denoted as $t$, to examine its impact on the performance of adversarial attacks. Due to space limitations, we report only the mean error growth rates (task-specific results are
cost and may alert the platform or API owner. Therefore, although such attacks may not yield optimal performance, conducting them under a strict query budget supports sustained adversarial behavior.
6 Conclusion

### v2

@@ IV-B LatentStealth
where $t$ denotes the iteration step, $\Delta{\hat{z}}^{(t)}$ represent the latent perturbation and its update at step $t$, $z$ is the latent representation of clean image, $(\alpha,\beta,\gamma)$ are the outputs obtained by feeding the clean image into $\mathcal{P}$, and $\xi_{z}$ is the learning rate for the gradient update in the latent space. The term $\frac{\partial\mathcal{G}_{\theta}}{\partial\Delta{\hat{z}}^{(t)}}$ denotes the Jacobian matrix of the decoder, whose computation is prohibitively expensive due to its high dimensionality ($n\times d$, often very large). Furthermore, each update necessitates a complete forward pass through the decoder network, resulting in substantial memory and computational consumption. Consequently, performing iterative updates in the latent space is highly inefficient (see Table V).
@@
TABLE I: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$)
@@
LatentStealth (Ours) 56.59 (37.89%) 9.74 (24.08%) 3.52 (18.92%) 128.57 (37.05%) 54.02 (51.74%) 41.83 (38.05%) 71.67 (43.92%) 9.93 (24.59%)
$\Delta x^{(0)}=x^{(0)}-x,\quad x^{(0)}=\mathcal{G}_{\theta}(z+\eta\cdot\epsilon_{2}).$ (17)
$\mathcal{P}(x+\Delta x^{(t)})=(\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),$ (18)
$\begin{split}\Delta x^{(t+1)}&=\Delta x^{(t)}+\\ &\quad\xi_{\Delta x}\cdot\nabla_{\Delta x^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma)),\end{split}$ (19)
$\nabla_{\Delta x^{(t)}}\mathcal{L}((\hat{\alpha}^{(t)},\hat{\beta}^{(t)},\hat{\gamma}^{(t)}),(\alpha,\beta,\gamma))=\frac{\partial\mathcal{L}}{\partial(x+\Delta x^{(t)})},$ (20)
where $\xi_{\Delta x}$ denotes the learning rate for gradient updates in pixel space. Importantly, computing gradients in pixel space does not require backpropagation through the decoder network. Instead, perturbations are applied directly to the original image $x$, thereby eliminating the need to compute the high-dimensional Jacobian transformation from latent to pixel space. This approach substantially reduces both the computational complexity and memory usage of each gradient update step.
$\mathcal{L}$ $=-\mathbb{E}_{\Delta x}\Big[\|\hat{\alpha}^{(t)}-\alpha\|_{2}^{2}+\|\hat{\beta}^{(t)}-\beta\|_{2}^{2}+\|\hat{\gamma}^{(t)}-\gamma\|_{2}^{2}\Big]$ (21)
$\quad+\frac{1}{w\times h}\sum_{i=0}^{w-1}\sum_{j=0}^{h-1}\big\|\Delta x^{(t)}(i,j)\big\|_{2}^{2}.$
where $w$ and $h$ denote the width and height of $x$, respectively. The first term encourages the perturbed pose parameters to diverge from those of $\mathcal{P}(x)$, thereby enhancing the effectiveness of the attack. The second term imposes a strict constraint on pixel-level variations to ensure the imperceptibility of the perturbation, effectively serving as a regularization term. According the Eqs 18, 20 and 21, the update rule for the pixel-space perturbation is formulated as:
$\Delta x^{(t+1)}=\Delta x^{(t)}+\xi_{\Delta x}\cdot(\frac{\partial\mathcal{L}}{\partial(x+\Delta x^{(t)})}-\lambda\cdot x^{(t)}),$ (22)
where $\lambda$ is a hyperparameter that balances attack strength and imperceptibility. It is evident that the computational complexity of updates in latent space is $\mathcal{O}(n\times d)$, whereas the complexity of pixel-space updates is only $\mathcal{O}(n)$. Therefore, the proposed noise enhancement strategy in pixel space incurs significantly lower computational overhead (see the Method C and LatentStealth in Table V). To provide a clearer understanding of the LatentStealth pipeline, we present a detailed the step of generating adversarial samples; see Algorithm 1.
Algorithm 1 Generation of Adversarial Sample
Input: Original Image $x$, Random Gaussian Noise $\epsilon_{1}$, $\epsilon_{2}$
Output: Adversarial Sample $\hat{x}$
Setting: Encoder of VAE $\mathcal{E}_{\phi}$, Decoder of VAE $\mathcal{G}_{\theta}$, EHPS model $\mathcal{P}$, Mean value $\mu_{\phi}$, Variance $\sigma_{\phi}^{2}$, Number of iterative queries $N$, Step size $\upsilon$, Noise magnitude $\eta$, Learning rate $\zeta$, Latent representation $z$, Multi-task loss function $\mathcal{L}$
1: Loading parameters $\phi$, $\theta$ of VAE
[…]
7:for $i$ = 1 to NN do
8:  $\hat{x}=x+\Delta x$
[…]
10:  $\mathcal{L}=\mathcal{L}(\hat{\alpha},\hat{\beta},\hat{\gamma},\hat{x},\alpha,\beta,\gamma,x)$
11:  Update
12:  $\Delta x=\Delta x-\zeta\nabla_{\Delta x}\mathcal{L}$
13:end for
14:return $\hat{x}=\Delta x+x$
V Experiments
V-A Experimental Settings
Datasets and EHPS Models. We evaluate our method on two benchmarks: 3DPW [35], with in-the-wild RGB videos and 3D SMPL pose annotations, and UBody [23], a large-scale upper-body mesh dataset with 1.05 million annotated images across 15 scenarios. For evaluation, we test LatentStealth on the three most representative EHPS models, including Hand4Whole [26], OSX [23], and SMPLer-X [2], which provides four ViT-based pre-trained variants, denoted as SMPLer-X-M, where M denotes the size of the Vision Transformer (ViT) backbone [1] (M $\in$ S, B, L, H).
TABLE II: Robustness evaluation of the state-of-the-art EHPS model SMPLer-X-H on the UBody dataset under adversarial perturbations from four types of random noise and a central white patch.
Attack PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$)
All Hands Face All Hands Face Body Hands
Clean 24.64 8.47 2.24 41.22 30.88 16.62 29.29 8.64
Random Noise $\uparrow$ 24.66 8.47 2.24 41.26 30.88 16.62 29.31 8.64
Random Noise $\uparrow\uparrow$ 24.86 8.44 2.24 41.98 31.06 16.76 29.57 8.61
Random Noise $\uparrow\uparrow\uparrow$ 25.34 8.43 2.25 44.29 31.44 17.07 30.22 8.59
Random Noise $\uparrow\uparrow\uparrow\uparrow$ 25.90 8.42 2.27 47.19 31.81 17.52 30.99 8.58
Patch 24.63 8.48 2.24 41.20 30.86 16.61 29.28 8.65
LatentStealth (Ours) 43.94 9.90 2.80 88.60 48.69 28.38 50.48 10.08
@@ V-A Experimental Settings
TABLE III: Performance comparison of different adversarial attack methods on state-of-the-art EHPS models on the 3DPW dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Attack MPJPE (Body) $\downarrow$ ($m$) PA MPJPE (Body) $\downarrow$ ($m$)
@@
LatentStealth (Ours) 128.72 (27.88%) 80.10 (17.92%)
Baseline and Evaluation Metrics. Currently, only one adversarial attack targets EHPS. To evaluate LatentStealth, we compare it with state-of-the-art methods in computer vision, including classical attacks (FGSM [7], PGD [25]), diffusion-based approaches (ACA [5], DiffAttack [4]), and the EHPS-specific TBA [22]. We report MPJPE and MPVPE for 3D joint and mesh accuracy, and PA-MPJPE/PA-MPVPE for alignment-aware evaluation, following prior EHPS benchmarks [26, 23, 2]. All metrics are in millimeters (mm).
Implementation Details. The experiments are conducted on a system running Ubuntu 20.04.6 LTS, using PyTorch framework, with a single NVIDIA L20 GPU. The VAE model employed is stabilityai/sd-vae-ft-mse from Hugging Face  [^1] An open source platform. https://huggingface.co/stabilityai/sd-vae-ft-mse. The maximum perturbation magnitude is set to $\eta=0.05$ with a step size of 0.01, and an $l_{\infty}$ constraint was enforced to ensure that changes in individual pixel values did not exceed 0.05. For baseline methods involving iterative optimization, the number of iterations is fixed at 20. For LatentStealth, the number of queries is set to $t=3$, and the balancing hyperparameter is fixed at $\lambda=0.5$.
V-B Robustness of EHPS Models
To evaluate the robustness of the EHPS model, we attack the state-of-the-art SMPLer-X-H using adversarial samples generated from four types of random noise and a central white patch on the UBody dataset. As shown in Table II, the performance metrics under these attacks deviate negligibly ($\leq$ 1%) from those on clean samples, indicating that ordinary noise does not meaningfully degrade EHPS model performance due to robust generalization from large, diverse training data. In contrast, our proposed LatentStealth successfully compromises SMPLer-X-H, revealing significant vulnerabilities to strategically crafted adversarial samples.
V-C Attack Performance on EHPS Models
TABLE IV: Quantitative comparison on UBody. The best result is highlighted.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ $_{2} \downarrow LL_{\infty}$ $\downarrow$
FGSM 26.91 0.45 0.48 52.01 34.72 0.10
PGD 24.90 0.33 0.48 48.96 43.72 0.10
ACA 26.99 0.92 0.21 18.18 35.48 0.11
DiffAttack 40.72 0.91 0.19 7.19 7.07 0.05
TBA 13.83 0.62 0.54 62.04 157.52 0.48
LatentStealth (Ours) 47.15 0.98 0.10 1.89 3.37 0.01
We first evaluate the effectiveness of LatentStealth on the UBody dataset, which includes full-body, hand, and facial regions of digital humans. As shown in Table I, across all tasks, the adversarial examples generated by LatentStealth cause the most significant degradation in the accuracy of all EHPS models, compared to other attack methods. Specifically, the average increase in estimation error ranged from 25.13% to 58.21%. Significantly, for the SMPLer-X-H model, the MPVPE (All) (sixth column) increased from 41.22 mm to 88.60 mm, corresponding to a 114.94% error growth rate.
TABLE V: Ablation study of different adversarial attack frameworks on state-of-the-art EHPS models on the UBody dataset. The error growth rates are marked in gray. The maximum error and maximum error growth rate on each setting are highlighted underlined.
Model Method PA MPVPE $\downarrow$ ($m$) MPVPE $\downarrow$ ($m$) PA MPJPE $\downarrow$ ($m$) Memory
@@ C Out of Memory $\geq$ 45 GB
LatentStealth (Ours) 50.16 (55.39%) 10.53 (4.78%) 3.33 (17.67%) 98.24 (71.66%) 59.06 (37.99%) 34.70 (55.88%) 58.62 (56.82%) 10.78 (4.97%) 7.47 GB
(a) UBody
(b) PW3D
Figure 5: Performance of different perturbation magnitude $\eta$.
(a) UBody
(b) PW3D
Figure 6: Performance of different test times $t$.
We further evaluate LatentStealth on the 3DPW dataset, focusing on how adversarial examples impact the estimation of body parts across different EHPS models. As shown in Table III, LatentStealth consistently causes the greatest reduction in model accuracy. Especially, LatentStealth yields a minimum error growth rate of 17.92% (increasing PA MPJPE from 67.93 mm to 80.10 mm on the Hand4Whole model) and a maximum of 48.46% (raising MPJPE from 79.46 mm to 118.00 mm on the SMPLer-X-B model), with average increases ranging from 17.27% to 44.49%. These results demonstrate that LatentStealth is highly destructive. Remarkably, its computational cost remains low, ensuring efficiency. Interestingly, we observe negative error growth rates in certain cases, such as PA MPVPE (Face) under the OSX model on UBody and several other metrics under specific settings. This finding suggests that while adversarial attacks generally impair EHPS performance, they may, in rare instances, also enhance estimation accuracy.
To further illustrate the effectiveness of LatentStealth, we conduct qualitative evaluations by visualizing the digital humans generated from both clean and adversarial samples produced by various attack methods. As shown in Figs. 3 and 4, LatentStealth leads to noticeable degradation in pose and shape estimation accuracy, resulting in significant deviations in the synthesized digital humans. These visual results clearly highlight the adverse impact of LatentStealth on the EHPS models’ ability to accurately reconstruct human poses.
V-D Stealthiness
We adopt six standard image quality evaluation metrics, e.g., PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index) [38], LPIPS (Learned Perceptual Image Patch Similarity) [40], FID (Fréchet Inception Distance) [9], $L_{2}$ error, and $L_{\infty}$ error, to assess the imperceptibility of the adversarial examples generated by our method. As presented in Table IV, our method achieves the best results in all metrics. Furthermore, we conduct additional qualitative experiments by visualizing the generated digital humans from both clean and adversarial samples generated by various attack methods across various scenarios to substantiate our findings. As shown in Figs. 3 and 4, LatentStealth causes a marked degradation in human pose and shape estimation accuracy, leading to evident distortions in the synthesized digital humans, and in some instances, rendering the generated poses entirely implausible. These results clearly demonstrate the adverse impact of LatentStealth on the ability of EHPS models to accurately capture human pose.
V-E Ablation Study
@@
Effects of Different Noise Injection Strategies. To further assess LatentStealth, we compare three alternative noise injection strategies within its framework: A) adding random noise in the pixel space and updating it iteratively via EHPS test; B) injecting random noise into the latent space after VAE encoding, then resampling and decoding to produce adversarial examples; C) injecting random noise into the VAE latent space, followed by resampling, decoding, and iterative updates via EHPS queries. We evaluate these strategies on SMPLer-X using the UBody dataset and report GPU memory usage. As shown in Table V, LatentStealth achieves strong attack performance with only 7.47–12.83 GB of memory, demonstrating its effectiveness and efficiency without requiring extensive resources.
@@
Effects of Different Perturbation Magnitude $\eta$. To examine the effect of perturbation magnitude $\eta$ on EHPS models, we evaluate SMPLer-X and report mean error growth rates on the UBody and PW3D datasets. As shown in Fig. 5, different values of $\eta$ yield similar attack effectiveness. This indicates that the model’s sensitivity to latent-space perturbations is concentrated along specific directions, rather than determined solely by magnitude. Consequently, increasing $\eta$ does not substantially enhance attack performance.
Effects of Different Test Times $t$. Similarly, we conduct an ablation study on the number of tests $t$ to the EHPS model, to evaluate its effect on attack performance. Due to space limits, we report only mean error growth rates. As shown in Fig. 6, $t$ is positively correlated with error growth, indicating that more tests enhance attack effectiveness. When $t=9$, the mean error growth rates for SMPLer-X-B, SMPLer-X-L, and SMPLer-X-H all exceed 100%. However, excessive tests raise computational cost and reduce efficiency. Although LatentStealth may not perform optimally at $t=3$, a low-cost attack remains attractive to adversaries.
VI Conclusion